# Can LLM-based Financial Investing Strategies Outperform the Market in Long Run?

### v4

@@ 6.3. Statistical Validation and Behavioural Diagnostics of LLM Agents
To validate our findings from the composite backtests and diagnose the underlying drivers of LLM agent performance, we conduct a unified statistical and behavioural analysis. First, we conduct paired t-tests comparing Buy and Hold, FinMem, and FinAgent across both Selected 4 (Table 3) and Composite (Table 4) setups. Second, we dissect the agents’ behavioural characteristics by examining their drawdown profiles, alpha ($\alpha$) and beta ($\beta$) decomposition, and trading turnover across the different selection environments. These metrics are obtained by regressing the strategy’s excess returns against the market’s excess returns based on the Capital Asset Pricing Model (CAPM) (capm-1964). The model is defined as: $R_{s}-R_{f}=\alpha+\beta(R_{m}-R_{f})+\epsilon$, where $R_{s}$ is the return of the strategy, $R_{m}$ is the market return, $R_{f}$ is the risk-free rate, and $\epsilon$ is the idiosyncratic residual. In this model, $\beta$ measures the strategy’s systematic risk or volatility relative to the market, while $\alpha$ represents the portion of the return not explained by market exposure, often considered a measure of strategy-specific skill.
@@
Table 5. Paired t-test p-values comparing Buy and Hold (B&H), FinMem, and FinAgent under Selected 4 and Composite setups.
Table 5 reports t-tests and p-values for the previous results, testing the null hypothesis of equal performance distributions. Under the selective period, statistical significance is inconsistent and limited mostly to individual stocks. However, after mitigating biases through the composite setup, the p-values drop substantially, indicating the market baseline (Buy and Hold) significantly outperforms both LLM strategies across all robust setups. Notably, while FinAgent tends to outperform FinMem when biases are controlled, both still significantly underperform simple market baselines. Furthermore, the behavioural analysis in Table 6 reveals that this underperformance is rooted in a lack of genuine skill; neither LLM agent generates statistically significant alpha, with all measured p-values exceeding 0.34. This finding robustly supports our main thesis that the claimed superiority of these models does not hold under rigorous evaluation, aligning with the Efficient Market Hypothesis (adaptive-markets-hypo).
A clear behavioural hierarchy also emerges between the two agents. FinMem consistently exhibits a more pathological trading profile, characterised by hyperactive turnover and poor risk management. Its commission ratio is approximately 5 to 9 times higher than FinAgent’s across both contexts, and its drawdown durations are substantially longer. This excessive, costly trading correlates with consistent value destruction, as evidenced by FinMem’s negative alpha in all scenarios. In contrast, FinAgent demonstrates a more controlled, albeit still unskilled, approach. Appendix F provides a comparative analysis with visualisations to further highlight the behavioural differences between FinMem and FinAgent.
These behaviours are directly modulated by the selection strategy, which acts as a powerful environmental filter. The Momentum selection strategy elicits the most engaged market posture from the agents, prompting their highest $\beta$ values. FinMem’s performance improves in this context relative to other environments, but it still yields a negative alpha of -1.34%. This is the only scenario where FinAgent produces a large positive alpha of +6.57%. Although this result lacks statistical significance (p=0.35), it suggests that the LLMs’ primary strength may not be in discovering novel signals but rather in exploiting strong, pre-existing market trends. In contrast, the Low Volatility environment prompts a risk-averse posture. Here, FinMem remains ineffective with a -1.04% alpha and a very low $\beta$ of 0.20. FinAgent also becomes highly conservative, with its risk profile improving (e.g., its average drawdown duration falls to 38.71 days) but at the cost of performance, generating a negative alpha.
@@ 7. Market Regime Analysis
Another key question in evaluating LLM-based investing strategies is whether they adapt appropriately across varying market conditions. Financial markets exhibit time-varying predictability and uncertainty across different economic, financial, and political regimes (kim2011stock). Some strategies may exploit these variations, while others may struggle to adapt. Distinct market environments—bull, bear, and sideways—present unique challenges and opportunities: bull markets reward aggressive positioning and high exposure, bear markets require effective risk management, and sideways markets test a strategy’s ability to navigate uncertainty in the absence of clear trends. By decomposing performance across these regimes, it is possible to determine whether strategies are overly conservative and miss opportunities during bullish periods, or excessively aggressive and incur significant losses during downturns. Understanding these regime-specific behaviours is essential for interpreting the strengths and weaknesses of LLM-based investing strategies and for guiding their future development (optimaltradingstrategyduringbullbear).
We label each calendar year based on the annual return of the S&P 500: $R_{y}=\frac{P_{T}-P_{0}}{P_{0}}$, where $P_{0}$ and $P_{T}$ are the adjusted closing prices on the first and last trading days of year $y$. A year is classified as bull if $R_{y}\geq+20\%$, bear if $R_{y}\leq-20\%$, and sideways otherwise. The $\pm 20\%$ threshold follows standard industry convention (wallstreetbullbear).
@@
These results suggest two directions for future LLM investors. A first direction concerns improving trend-detection so the strategy can at least match passive equity beta in up-markets. A second direction relates to embedding explicit regime-aware risk controls that scale exposure down as volatility or draw-down risk rises. Balancing risk-taking and risk management, aggression and defence, rather than increasing model size, appears the key to closing the gap with traditional methods.
@@ 8. Findings and Takeaways
First, we find that LLM-derived alpha is likely a methodological artefact of narrow, biased evaluations. The performance advantages reported in short-term, selective studies vanish under our bias-mitigated backtests, which reveal a consistent and statistically significant failure to generate alpha (Section 6.3). This suggests that current LLMs do not overcome the Efficient Market Hypothesis (fama1970efficient) in realistic conditions, and that prior gains stemmed from survivorship and look-ahead biases rather than genuine market inefficiency.
Second, model complexity does not equate to market competence. The scaling laws of natural language processing (kaplan2020scalinglawsneurallanguage) do not translate effectively to financial markets, which impose intrinsic limits on extractable signals (backtesting-ssrn). We show that larger models do not reliably outperform smaller ones, and both are consistently bettered by simpler models like ARIMA on risk-adjusted metrics (Table 4). Without encoded financial logic, architectural complexity appears to add noise rather than value.
Third, we diagnose “how” LLM agents fail, revealing a fundamental misalignment with market regimes. Our analysis (§7, Appendix F) shows that agents are pathologically miscalibrated: they are too conservative in bull markets and too aggressive in bear markets. This behavioural flaw contradicts the Adaptive Markets Hypothesis (adaptive-markets-hypo), shifting the issue from merely a lack of profitability to a more profound failure in the agents’ decision-making policies.
@@ Appendix A Data Collection
We collect daily price data for over 7,000 U.S. equities spanning from 2000 to 2024. In addition to actively traded stocks, our dataset includes delisted symbols that were historically part of the S&P 500 index, based on the archived constituent list [^5] https://github.com/fja05680/sp500. This inclusion enhances the historical completeness of our dataset and mitigates survivorship bias within the context of index-based evaluations.
@@
The financial news dataset, initially compiled by 10.1145/3637528.3671629, comprises 15.7 million records pertaining to 4,775 S&P 500 companies, spanning the years 1999 to 2023. We have organised the news by aligning it with the respective companies and indexing it by date.
@@ B.1. Timing-based Strategies
This category includes FinMem (yu2023finmemperformanceenhancedllmtrading) and FinRobot (yang2024finrobotopensourceaiagent). We acknowledge other works, such as FinCon (yu2024fincon) and MarketSenseAI (fatouros2025marketsenseai20enhancingstock), but they are not (yet) open-source, which prevents us from generating backtesting results.
@@
We implement and cover several well-known traditional rule-based (indicator-based) investing strategies, such as Buy and Hold, Simple Moving Average Crossover, Weighted Moving Average Crossover, ATR Band, Bollinger Bands (bollinger2002bollinger), Trend Following (Wilcox2009trendfollowing), and Turn of the Month (turnofthemonth2008). These strategies typically rely on one or multiple technical indicators or domain-based rules to generate timely buy/sell signals, aiming to exploit identifiable market patterns or anomalies.
@@
In contrast to fixed rules or indicator-based triggers, these strategies rely on data-driven models—often statistical or neural network forecasters—to predict future price movements. Specifically, they buy or hold if an uptrend is indicated and sell (or go short) otherwise. This can be viewed as a relatively naive application of ML/DL forecasters, but it is widely used as a benchmark method for such models. Although one could consider the forecast output as a type of “indicator”, the reliance on predictive algorithms capable of uncovering complex patterns sets these methods apart from purely rule-based approaches. We include the well-known ARIMA (10.5555/574978) and XGBoost (10.1145/2939672.2939785) in this category and also cover forecasters based on LLMs, differentiating them from the LLM investors.
@@
We also implement widely used RL algorithms for financial markets, including Advantage Actor-Critic (A2C), Proximal Policy Optimisation (PPO), Twin Delayed Deep Deterministic Policy Gradient (TD3), and Soft Actor-Critic (SAC), utilising the FinRL framework (finrl-2022; finrl-2024). Each agent learns investing policies by interacting with a simulated trading environment based on the OpenAI Gym API, using real historical market data.
  • Environment design. The trading process is modelled as a Markov Decision Process (MDP), where the agent observes a market state, selects an action, and receives a reward at each step. The environment simulates live stock markets and follows the time-driven simulation principle.
  • State space. At each timestep, the agent receives a state vector capturing account balance, current shares held for each asset, stock prices (close, open, high, low), trading volumes, and technical indicators (e.g., MACD, RSI). The feature set is configurable and can be aggregated at daily, hourly, or minute frequency.
  • Action space. The action space allows buying, holding, or selling any subset of assets. Typically, actions are discrete values representing the number of shares to transact per asset (e.g., $a\in\{-k,\ldots,-1,0,1,\ldots,k\}$ for each stock), where negative, zero, and positive values denote selling, holding, and buying, respectively.
  • Reward function. The primary reward is the change in portfolio value after each trading action, i.e., $r(s,a,s^{\prime})=v^{\prime}-v$, where $v$ and $v^{\prime}$ are the portfolio values at states $s$ and $s^{\prime}$.
@@ B.2. Selection-based Strategies
Following the well-documented momentum factor (Muller31032010), this strategy selects the stocks with the highest recent price appreciation. For each candidate stock, we calculate a momentum score based on its historical price data. Specifically, the score is the percentage return over a “momentum period” (e.g., 100 trading days), but we exclude the most recent “skip period” (e.g., 21 trading days) from the calculation. This practice is common in momentum strategies to avoid the “short-term reversal” effect (on-persistence-1997). The score for a given stock is calculated as shown in Equation 1. $t$ is the selection date. All candidate stocks are then ranked in descending order by this score, and the top-$k$ stocks (e.g., $k=5$) are selected.
(1) $\text{Momentum Score}=\frac{\text{Price}_{t-\text{skip\_period}}}{\text{Price}_{t-\text{momentum\_period}}}-1$
@@
This strategy is based on the “volatility effect” anomaly, where low-volatility stocks have been empirically shown to generate higher risk-adjusted returns (volatility-effect). For each candidate stock, we measure its historical volatility over a recent “look-back period” (e.g., 21 trading days). The volatility is calculated as the standard deviation of its weekly log returns within this period. We use weekly returns ($\ln(P_{t}/P_{t-5})$) rather than daily returns to smooth out daily noise. Candidate stocks are then ranked in ascending order by their calculated volatility, and the top-$k$ stocks with the lowest volatility are selected for the portfolio.
@@
Unlike the single-factor methods above, the FinCon Selection Agent (yu2024fincon) aims to construct a diversified portfolio by explicitly considering both performance and inter-stock correlation. Its selection process is more sophisticated:
@@
    Primary Selection: The agent first ranks each stock using a combined score that balances risk-adjusted return (Sharpe ratio) and its potential for diversification (low average correlation with all other stocks, $\overline{\rho}$). The score is calculated as:
    $\text{Score}=\text{Sharpe Ratio}\times(1-\overline{\rho})$
    The top-$k$ stocks based on this score form the initial portfolio.
@@
This two-stage process ensures that the final selection prioritises a balance of individual stock quality and overall portfolio diversification.
@@ Appendix D Extra Results on Selective Symbols
Tables 2 and 7 further substantiate our findings by highlighting the performance instability of FinMem and FinAgent when extending evaluation periods even marginally. Specifically, extending the evaluation by just two months beyond the originally reported periods (yu2023finmemperformanceenhancedllmtrading) results in notable inconsistencies in critical performance metrics. It should be noted that the results for the LLM strategies are retrieved from yu2024fincon, while the traditional rule-based results presented are based on our implementations.
@@
Rule
Based
Turn of The Month 0.928 27.850 -11.642 0.123 3.487 -14.892 0.874 31.344 -17.995 0.407 7.744 -11.955
@@
Rule
Based
Turn of The Month 0.098 3.337 -12.498 0.343 7.188 -13.519 0.987 18.942 -10.641 -0.020 -8.999 -33.895
@@
Table 7. Backtest performance of traditional rule-based (indicator-based) strategies and FinCon over the selective period (2022-10-05 to 2023-06-10), as presented in yu2024fincon, evaluated using four metrics: cumulative return (CR), Sharpe ratio (SPR), annual volatility (AV), and maximum drawdown (MDD). The best metrics are highlighted in red, while the second best are marked in blue. “-” metrics across the board indicate no trade signals were triggered.

### v5

@@ 6.3. Statistical Validation and Behavioural Diagnostics of LLM Agents
To validate our findings from the composite backtests and diagnose the underlying drivers of LLM agent performance, we conduct a unified statistical and behavioural analysis. First, we conduct paired t-tests comparing Buy and Hold, FinMem, and FinAgent across both Selected 4 (Table 3) and Composite (Table 4) setups. Second, we dissect the agents’ behavioural characteristics by examining their drawdown profiles, alpha ($\alpha$) and beta ($\beta$) decomposition, and trading turnover across the different selection environments. These metrics are obtained by regressing the strategy’s excess returns against the market’s excess returns based on the Capital Asset Pricing Model (CAPM) (Sharpe, 1964). The model is defined as: $R_{s}-R_{f}=\alpha+\beta(R_{m}-R_{f})+\epsilon$, where $R_{s}$ is the return of the strategy, $R_{m}$ is the market return, $R_{f}$ is the risk-free rate, and $\epsilon$ is the idiosyncratic residual. In this model, $\beta$ measures the strategy’s systematic risk or volatility relative to the market, while $\alpha$ represents the portion of the return not explained by market exposure, often considered a measure of strategy-specific skill.
@@
Table 5. Paired t-test p-values comparing B&H, FinMem, and FinAgent under Selected 4 and Composite setups.
Table 5 reports t-tests and p-values for the previous results, testing the null hypothesis of equal performance distributions. Under the selective period, statistical significance is inconsistent and limited mostly to individual stocks. However, after mitigating biases through the composite setup, the p-values drop substantially, indicating the market baseline (B&H) significantly outperforms both LLM strategies across all robust setups. Notably, while FinAgent tends to outperform FinMem when biases are controlled, both still underperform simple market baselines. Furthermore, the behavioural analysis in Table 6 reveals that this underperformance is rooted in a lack of genuine skill; neither LLM agent generates statistically significant alpha, with all measured p-values exceeding 0.34. This finding robustly supports our main thesis that the claimed superiority of these models does not hold under rigorous evaluation, aligning with the Efficient Market Hypothesis (Lo, 2004).
A clear behavioural hierarchy emerges between the two agents. FinMem consistently shows a more pathological trading profile, marked by excessive turnover and poor risk management. Its commission ratio is five to nine times higher than FinAgent’s across both contexts, and its drawdown durations are substantially longer. This overtrading leads to persistent value destruction, reflected in FinMem’s negative alpha in all scenarios. In contrast, FinAgent follows a more restrained, though still unskilled, trading strategy. Appendix F provides a comparative analysis with visualisations to further highlight the behavioural differences between FinMem and FinAgent as supplementary evidence.
These behaviours are directly modulated by the selection strategy, which acts as a powerful environmental filter. The Momentum selection strategy elicits the most engaged market posture from the agents, prompting their highest $\beta$ values. FinMem’s performance improves in this context relative to other environments, but it still yields a negative alpha of -1.34%. This is the only scenario where FinAgent produces a large positive alpha of +6.57%. Although this result lacks statistical significance (p=0.35), it suggests that the LLMs’ primary strength may not be in discovering novel signals but rather in exploiting strong, pre-existing market trends. In contrast, the Low Volatility environment takes a risk-averse posture. Here, FinMem remains ineffective with a -1.04% alpha and a very low $\beta$ of 0.20. FinAgent also becomes highly conservative, with its risk profile improving (e.g., its average drawdown duration falls to 38.71 days) but at the cost of performance, generating a negative alpha.
@@ 7. Market Regime Analysis
Another key question in evaluating LLM-based investing strategies is whether they adapt appropriately across varying market conditions. Financial markets exhibit time-varying predictability and uncertainty across different economic, and political regimes (Kim et al., 2011). Some strategies may exploit these variations, while others may struggle to adapt. Distinct market environments—bull, bear, and sideways—present unique challenges and opportunities: bull markets reward aggressive positioning and high exposure, bear markets require effective risk management, and sideways markets test a strategy’s ability to navigate uncertainty in the absence of clear trends. By decomposing performance across these regimes, it is possible to determine whether strategies are overly conservative and miss opportunities during bullish periods, or excessively aggressive and incur significant losses during downturns. Understanding these regime-specific behaviours is essential for interpreting the strengths and weaknesses of LLM-based investing strategies (Hui and Chan, 2018).
We label each calendar year based on the annual return of the S&P 500: $R_{y}=\frac{P_{T}-P_{0}}{P_{0}}$, where $P_{0}$ and $P_{T}$ are the adjusted closing prices on the first and last trading days of year $y$. A year is classified as bull if $R_{y}\geq+20\%$, bear if $R_{y}\leq-20\%$, and sideways otherwise. The $\pm 20\%$ threshold follows standard industry convention (Zweig, 2019).
@@
These results suggest two directions for future LLM investors. First, trend-detection capabilities to ensure that the strategy can at least match passive equity beta during upward market phases. Second, incorporating explicit regime-aware risk controls that reduce exposure as volatility or drawdown risk increases. Balancing risk-taking and risk management, rather than simply increasing model size, appears the key to closing the gap with traditional methods.
@@ 8. Findings and Takeaways
First, we find that LLM-derived alpha is likely a methodological artefact of narrow, biased evaluations. The performance advantages reported in short-term, selective studies vanish under our bias-mitigated backtests, which reveal a consistent and statistically significant failure to generate alpha (§6.3). This suggests that current LLMs do not overcome the Efficient Market Hypothesis (Fama, 1970) in reality, and that prior gains stemmed from survivorship and look-ahead biases rather than genuine market inefficiency.
Second, model complexity does not equate to market competence. The scaling laws of natural language processing (Kaplan et al., 2020) do not translate effectively to financial markets, which impose intrinsic limits on extractable signals (Harvey and Liu, 2013). We show that larger models do not reliably outperform smaller ones, and both are consistently bettered by simpler models like ARIMA on risk-adjusted metrics (Table 4). Without encoded financial logic, architectural complexity appears to add noise rather than value.
Third, we diagnose “how” LLM agents fail, revealing a fundamental misalignment with market regimes. Our further analysis (§7, Appendix F) shows that agents are pathologically miscalibrated: they are too conservative in bull markets and too aggressive in bear markets. This behavioural flaw contradicts the Adaptive Markets Hypothesis (Lo, 2004), shifting the issue from merely a lack of profitability to a more profound failure in the agents’ decision-making policies.
@@ 9. Conclusion
A remaining limitation is potential data leakage, as some evaluation data may have been included in the pretraining corpora of proprietary LLMs and cannot be fully verified. However, any such leakage would bias results in favour of LLMs and therefore does not alter our central findings.
@@
Limitations
There are several limitations to our current study. First, we did not individually tune the traditional rule-based strategies for each rolling evaluation window. Typically, applying domain-specific market insights to optimise parameters can significantly enhance the performance of these methods. However, we argue that our current configuration remains valid and effectively demonstrates the competitive disadvantage faced by LLM strategies. Indeed, tuning the parameters of traditional rule-based strategies would likely elevate their performance further, reinforcing rather than undermining our main conclusions.
Second, our evaluation has not fully eliminated look-ahead bias. Pre-trained LLMs, due to their inherent training corpus, may inadvertently contain stock-related information from historical periods overlapping our test sets. Despite this potential data leakage, the observed underperformance of LLM strategies strengthens our critical assessment. Explicitly addressing this look-ahead concern through controlled model training or careful exclusion of financial data from training corpora will be an important avenue for future research.
Third, to ensure experiment reproducibility, we restricted our analysis to publicly available data, excluding proprietary sources such as private newsfeeds, earning transcripts, or expert analyses. Nonetheless, the FINSABER framework was deliberately designed to be modular and extensible, allowing researchers with access to private data to easily integrate additional information sources. Our primary goal remains providing a rigorous, long-term evaluation pipeline that minimises selective reporting. Researchers lacking proprietary data can fully replicate our results using openly accessible resources.
Acknowledgements
We thank the reviewers and the area chair for their useful feedback. The authors acknowledge the use of resources provided by the Edinburgh Compute and Data Facility [^5] http://www.ecdf.ed.ac.uk/ (ECDF).
@@ Appendix A Data Collection
We collect daily price data for over 7,000 U.S. equities spanning from 2000 to 2024. Additionally, our dataset includes delisted symbols that were historically part of the S&P 500 index, based on the archived constituent list. This inclusion enhances the historical completeness of our dataset and mitigates survivorship bias within the context of index-based evaluations.
@@
The financial news dataset, initially compiled by Dong et al. (2024), comprises 15.7 million records pertaining to 4,775 S&P 500 companies, spanning the years 1999 to 2023. We have organised the news by aligning it with the respective companies and indexing it by date.
@@ B.1. Timing-based Strategies
This category includes FinMem (Yu et al., 2023) and FinRobot (Yang et al., 2024). We acknowledge other works, such as FinCon (Yu et al., 2024) and MarketSenseAI (Fatouros et al., 2025), but they are not (yet) open-source, which prevents us from generating backtesting results.
@@
We implement and cover several well-known traditional rule-based (indicator-based) investing strategies, such as Buy and Hold, Simple Moving Average Crossover, Weighted Moving Average Crossover, ATR Band, Bollinger Bands (Bollinger, 2002), Trend Following (Wilcox et al., 2005), and Turn of the Month (McConnell and Xu, 2008). These strategies typically rely on one or multiple technical indicators or domain-based rules to generate timely buy/sell signals, aiming to exploit identifiable market patterns or anomalies.
@@
In contrast to fixed rules or indicator-based triggers, these strategies rely on data-driven models (statistical or neural network forecasters) to predict future price movements. Specifically, they buy or hold if an uptrend is indicated and sell (or go short) otherwise. This can be viewed as a relatively naive application of ML/DL forecasters, but it is widely used as a benchmark method for such models. Although one could consider the forecast output as a type of “indicator”, the reliance on predictive algorithms capable of uncovering complex patterns sets these methods apart from purely rule-based approaches. We include the well-known ARIMA (Box and Jenkins, 1990) and XGBoost (Chen and Guestrin, 2016) in this category and also cover forecasters based on LLMs, but these are not LLM investors.
@@
We also implement widely used RL algorithms for financial markets, including Advantage Actor-Critic (A2C), Proximal Policy Optimisation (PPO), Twin Delayed Deep Deterministic Policy Gradient (TD3), and Soft Actor-Critic (SAC), utilising the FinRL framework (Liu et al., 2022, 2024). Each agent learns investing policies by interacting with a simulated trading environment based on the OpenAI Gym API, using real historical market data.
@@ B.2. Selection-based Strategies
Following the well-documented momentum factor (Muller and Ward, 2010), this strategy selects the stocks with the highest recent price appreciation. For each candidate stock, we calculate a momentum score based on its historical price data. Specifically, the score is the percentage return over a “momentum period” (e.g., 100 trading days), but we exclude the most recent “skip period” (e.g., 21 trading days) from the calculation. This practice is common in momentum strategies to avoid the “short-term reversal” effect (Carhart, 1997). The score for a given stock is calculated as: $\text{Momentum Score}=(\text{Price}_{t-\text{skip\_period}})/(\text{Price}_{t-\text{momentum\_period}})-1$. $t$ is the selection date. All candidate stocks are then ranked in descending order by this score, and the top-$k$ stocks (e.g., $k=5$) are selected.
@@
This strategy is based on the “volatility effect” anomaly, where low-volatility stocks have been empirically shown to generate higher risk-adjusted returns (Blitz and Vliet, 2007). For each candidate stock, we measure its historical volatility over a recent “look-back period” (e.g., 21 trading days). The volatility is calculated as the standard deviation of its weekly log returns within this period. We use weekly returns ($\ln(P_{t}/P_{t-5})$) rather than daily returns to smooth out daily noise. Candidate stocks are then ranked in ascending order by their calculated volatility, and the top-$k$ stocks with the lowest volatility are selected for the portfolio.
@@
Unlike the single-factor methods above, the FinCon Selection Agent (Yu et al., 2024) aims to construct a diversified portfolio by explicitly considering both performance and inter-stock correlation. Its selection process is more sophisticated:
@@
    Primary Selection: The agent first ranks each stock using a combined score that balances risk-adjusted return (Sharpe ratio) and its potential for diversification (low average correlation with all other stocks, $\overline{\rho}$). The score is calculated as: $\text{Score}=\text{Sharpe Ratio}\times(1-\overline{\rho})$ The top-$k$ stocks based on this score form the initial portfolio.
@@ Appendix D Extra Results on Selective Symbols
Tables 2 and 7 further substantiate our findings by highlighting the performance instability of FinMem and FinAgent when extending evaluation periods even marginally. Specifically, extending the evaluation by just two months beyond the originally reported periods (Yu et al., 2023) results in notable inconsistencies in critical performance metrics. It should be noted that the results for the LLM strategies are retrieved from Yu et al. (2024), while the traditional rule-based results presented are based on our implementations.
@@
Rule Based Turn of The Month 0.928 27.850 -11.642 0.123 3.487 -14.892 0.874 31.344 -17.995 0.407 7.744 -11.955
@@
Rule Based Turn of The Month 0.098 3.337 -12.498 0.343 7.188 -13.519 0.987 18.942 -10.641 -0.020 -8.999 -33.895
@@
Table 7. Backtest performance of traditional rule-based (indicator-based) strategies and FinCon over the selective period (2022-10-05 to 2023-06-10), as presented in Yu et al. (2024), evaluated using four metrics: cumulative return (CR), Sharpe ratio (SPR), annual volatility (AV), and maximum drawdown (MDD). The best metrics are highlighted in red, while the second best are marked in blue. “-” metrics across the board indicate no trade signals were triggered.